# Containerization in Multi-Cloud Environment: Roles, Strategies, Challenges, and Solutions for Effective Implementation

### v2

@@ 4.6. Deployment Challenges and Solutions Framework (RQ4 - RQ5)
Standardization and Automation in Deployment: Standardization and automation stand out as foundational pillars for achieving streamlined and uniform deployments in multi-cloud landscapes. As Raj et al. (Raj et al., 2018) highlighted, embracing automated deployment pipelines paired with standardized toolsets can significantly diminish complexity and mitigate human-induced errors within multi-cloud frameworks. The challenges, captured under Performance Testing Deployment, Deployment Validation, and Cloud Management Platform Evaluation Deployment in the table, elucidate the nuanced complexities associated with deployments spanning multiple cloud platforms. Solutions such as Standardized Benchmarking and Automation, Testing Process Management System, and Standardized Output Formats and Evaluation Criteria emphasize the imperative nature of adopting uniform procedures and harnessing automated solutions. Multi-cloud deployments, inherently intricate, benefit profoundly from standardization, bringing forth predictability and coherence, while automation paves the way for diminishing human-driven inconsistencies and bolstering deployment speeds. Implications: For researchers, the quest for optimized frameworks championing standardization and automation within multi-cloud environment presents a promising research trajectory. For practitioners, adeptness with standardized methodologies and tools, combined with proficiency in automation implementation, becomes crucial in navigating the multi-cloud management maze efficiently
@@
Security and Compliance in multi-cloud environment: Security remains paramount in cloud deployments. Fernandes et al. underscore that the labyrinth of diverse platforms and differing standards in multi-cloud configurations amplifies security complexities (Fernandes et al., 2014). Challenges articulated as Security Deployment in Multi-container Environment and Malicious Service Management Deployment in Table LABEL:tab:Deployment-solutions accentuate the primacy of security considerations. Propounded solutions like Cross-Container Isolation and Collusion-Resilient Trust Aggregation Technique spotlight the industry’s shift towards advanced and meticulous security protocols within multi-cloud paradigms. The ever-evolving landscape of multi-cloud deployments demands security solutions that are not only rigorous but also malleable and formidable. To attain comprehensive security across diverse platforms, strategies must encapsulate both depth of understanding and breadth of application. Implications: For the academic community, pioneering avant-garde security solutions apt for multi-cloud environment offers a fertile domain of inquiry. Meanwhile, practitioners should enshrine principles of data integrity, trust assurance, and vulnerability mitigation at the heart of their multi-cloud deployment blackprints.
@@
Integration, Monitoring, and Infrastructure Management: As highlighted by Ferrer et al. (Ferrer et al., 2016) the ability to operate seamlessly across disparate cloud providers and maintain flexible infrastructures is paramount in harnessing the full potential of multi-cloud deployments. The presented challenges, including Network Latency Deployment and Vendor Lock-in Prevention Deployment, underscore the indispensability of agility and cross-platform compatibility. Proposed solutions such as VM Selection for Composite Applications and Interoperability Layer Above Cloud Infrastructure accentuate the importance of crafting adaptable infrastructure and formulating strategies to mitigate vendor entrenchment. In the context of multi-cloud environment, flexibility and interoperability transcend mere advantageous attributes; they become pivotal determinants shaping adaptability, scalability, and the sustained success of deployment blackprints. Implications: For the academic community, delving into novel approaches that champion infrastructure pliability and seamless operation across diverse cloud platforms emerges as a prospective research avenue. For industry professionals, anchoring strategies around flexibility and interoperability not only optimizes current deployments but also fortifies them against future technological evolutions and shifting business dynamics.
4.7. Monitoring Challenges and Solutions (RQ4 - RQ5)
@@
Enhancing Performance Stability through Multi-Cloud Monitoring: In the context of multi-cloud environment, ensuring consistent and reliable performance monitoring for containerized applications can be challenging due to the dynamic nature of cloud resources. Existing studies (e.g., (Zhong et al., 2022) (Mungoli, 2023)) highlighted the impact of varying resource availability on the performance of containerized applications across multiple clouds. They emphasized the need for performance benchmarking to track fluctuations and maintain a consistent user experience. The solutions presented in our framework, such as benchmarking performance variation and utilizing Kubernetes for monitoring, align with existing studies suggestion for benchmarking. Our proposed framework expands on this by incorporating Kubernetes for improved multi-cloud performance monitoring. Our framework also suggest reliable performance monitoring requires a combination of standardized benchmarking techniques and cloud-specific tools like Kubernetes. This approach can effectively address performance variations and ensure a consistent user experience across diverse cloud environments. Implications: For practitioners, adopt benchmarking practices to track performance fluctuations, and leverage Kubernetes for unified multi-cloud performance monitoring. For researchers, explore further refinements in benchmarking methodologies and investigate the impact of dynamic multi-cloud environment on performance consistency.
@@
Streamlining Resource and Network Management across Multi-Clouds: Managing network communications, system images, and resource utilization in a multi-cloud setting poses challenges due to the diversity of cloud platforms. Sedghpour et al. discussed the complexities of managing network communications in multi-cloud environment. They emphasized the need for standardized approaches like service mesh to enhance network visibility and control (Sedghpour and Townend, 2022). Solutions in our proposed framework, such as using Docker Server for containerized services and Prometheus/Grafana for resource monitoring, align with Sedghpour et al.’s (Sedghpour and Townend, 2022) suggestion for standardized approaches. Additionally, our framework extends these solutions to address broader resource management challenges. Employing standardized tools like Docker Server, along with platforms like Prometheus and Grafana, can help streamline resource management and network communication monitoring across diverse cloud platforms. Implications: For practitioners, implement Docker Server and utilize platforms like Prometheus and Grafana for efficient resource monitoring and management across multi-cloud environment. For researchers, investigate further integration possibilities for standardized resource management tools and explore the impact of such tools on overall system performance and reliability.
@@
Strengthening Security, Compliance, and Trust: Ensuring data privacy, continuous security control, and compliance in multi-cloud containerized applications is essential but challenging due to the distributed nature of cloud resources. Existing studies (e.g., (Alonso et al., 2023b) (Raj et al., 2018)) discussed the challenges of ensuring security and compliance across multi-cloud environment. They highlighted the importance of continuous monitoring and enforcement mechanisms to address security and compliance gaps. Solutions presented in our proposed framework, such as using machine learning for performance supervision and employing security assurance platforms, are aligned with existing studies that emphasis on continuous monitoring and enforcement for security and compliance. Implementing advanced techniques like machine learning for monitoring and adopting security assurance platforms can significantly enhance security, compliance, and trust management in multi-cloud containerized applications. Implications: For practitioners, integrate machine learning-based monitoring and adopt security assurance platforms to ensure continuous security control and compliance across diverse cloud environments. For researchers, explore the scalability and effectiveness of machine learning-driven monitoring techniques and investigate novel methods for enforcing security and compliance in multi-cloud setups.
4.8. Tools and Frameworks (RQ6)
@@
Diverse Cloud Services by Leading Providers: Our study reveals a diverse array of cloud services, including IaaS, PaaS, and SaaS offerings from major providers like AWS, Azure, and Google Cloud, emphasizing the extensive options available to developers (see Table 15). This underscores the significant range of choices available to developers. Noteworthy cloud service providers highlighted in our study include Amazon EC2, Google Cloud App Engine, and Azure, further underscoring this finding. A pivotal insight gleaned from our research underscores the paramount importance of harmonizing cloud service models with specific application prerequisites, thereby facilitating efficient allocation of resources and optimization of costs. Implications: For practitioners, our outcomes underscore the critical necessity of meticulously evaluating and selecting the most fitting service model based on the unique demands of a given project. This strategic approach ensures the best alignment between technology and goals. Moreover, the implications for researchers are promising: an avenue emerges for in-depth exploration of trade-offs and the identification of optimal practices while navigating the intricate array of cloud service models. This exploration holds potential for enhancing both the efficiency and effectiveness of cloud solutions.
@@
Docker and Kubernetes: Pioneers of Cross-Cloud Containerization: Our investigation spotlights the profound significance of Docker and Kubernetes in streamlining the containerization and orchestration of applications across a spectrum of cloud platforms. Docker adeptly encapsulates applications and their dependencies within self-contained containers, while Kubernetes takes center stage as the orchestration juggernaut, automating deployment, scaling, and management of these containers. These technologies receive direct validation through table references, serving as prime exemplars. Our research yields a compelling revelation: Docker and Kubernetes hold pivotal roles in ensuring consistent application deployment and seamless scalability, effectively simplifying the complexities of multi-cloud environment. Implications: Practitioners stand to gain significantly by adopting Docker and Kubernetes, as this potent combination bolsters efficiency and cultivates heightened portability. Meanwhile, researchers have a promising avenue to explore deeper facets, including security, performance optimization, and the intricate domains of advanced use cases for Docker and Kubernetes within multi-cloud landscapes. This exploration promises valuable insights that can finely calibrate the potential of these cutting-edge technologies.
@@
The Rise of DevOps: Embracing Automation Tools: Our exploration spotlights the important role of DevOps practices and tools such as Puppet, Chef, and Terraform within contemporary software development landscapes. This emphasis revolves around the power of automation in facilitating streamlined deployment and management processes. These tools, by enabling consistent provisioning and deployment of infrastructure, effectively curtail errors and ensure the replicability of processes. The table substantiates this notion through references to Puppet, Chef, and Terraform. Our study unearths a pivotal insight: DevOps practices wield the ability to augment the software development life cycle by automating tasks, ushering in a realm of efficient and dependable deployment. Implications: Practitioners are advised to seamlessly integrate DevOps principles into their workflows by harnessing the prowess of these tools. Furthermore, a realm of exploration opens for researchers, inviting deeper analyses into the evolutionary trajectory of DevOps practices and their profound impact on elevating software quality and hastening development velocities.
5. Threats to Validity
@@
This systematic mapping study is susceptible to various threats that could influence its outcomes. To address these potential threats, we adhered to the established guidelines for conducting SMSs and SLRs as outlined in (Keele et al., 2007) and (Petersen et al., 2008b). We further categorized and analyzed the validity threats to our study based on the four distinct types of validity threats mentioned in (Zhou et al., 2016) and (Wohlin et al., 2012). In the subsequent sections, we delve into the specific validity threats that pertain to the different phases of this SMS.
5.1. Internal Validity
@@
  • Study Selection: We have outlined the study screening and selection process in Table 3. To ensure objectivity and eliminate personal bias in study selection, we implemented a two-phase approach: (i) initial study screening, and (ii) qualitative evaluation of the shortlisted studies. During this procedure, the lead two authors conducted the study screening based on the criteria explicitly detailed in Section 3. To assess the objectivity of the screening process and the level of agreement between the two authors, we applied Cohen’s Kappa (Cohen, 1968). In cases where the lead authors did not agree, the second and third authors independently reviewed the disputed studies to reach a consensus. All the researchers involved in this study have extensive knowledge and research experience in containerization applications within a multi-cloud environment.
  • Data Extraction: Bias from researchers during data extraction can pose a significant challenge in both SMSs and SLRs. To address this potential issue, we developed a standardized data extraction form (see Table 5) to ensure consistent data retrieval. Initially, the first and fourth authors extracted the data. If any uncertainties arose regarding the extracted data, comprehensive discussions were convened among all authors. Following the recommendations in (Wohlin et al., 2012), a subset of the extracted data was cross-verified by the second and third authors.
  • Bias on Themes Classification: Misclassification of data and primary studies may result in biases stemming from subjective interpretation. To minimize this risk, we adhered to the thematic analysis guidelines established by Braun et al. (Braun and Clarke, 2006), and implemented a six-step process for thematic classification, as detailed in Section 2.4.
  • Data Synthesis: We employed both qualitative and quantitative approaches to assess the gathered data. Potential biases in data synthesis could influence the interpretation of our findings. To address this concern, we used open coding and constant comparison techniques from Grounded Theory (Glaser and Strauss, 1967) to analyze the qualitative data extracted from the selected studies.
5.2. External Validity

### v3

@@ 4.6 Deployment Challenges and Solutions Framework (RQ4 - RQ5)
Standardization and Automation in Deployment: Standardization and automation stand out as foundational pillars for achieving streamlined and uniform deployments in multi-cloud landscapes. As Raj et al. [6] highlighted, embracing automated deployment pipelines paired with standardized toolsets can significantly diminish complexity and mitigate human-induced errors within multi-cloud frameworks. The challenges, captured under Performance Testing Deployment, Deployment Validation, and Cloud Management Platform Evaluation Deployment in the table, elucidate the nuanced complexities associated with deployments spanning multiple cloud platforms. Solutions such as Standardized Benchmarking and Automation, Testing Process Management System, and Standardized Output Formats and Evaluation Criteria emphasize the imperative nature of adopting uniform procedures and harnessing automated solutions. Multi-cloud deployments, inherently intricate, benefit profoundly from standardization, bringing forth predictability and coherence, while automation paves the way for diminishing human-driven inconsistencies and bolstering deployment speeds. Multi-cloud deployments, inherently intricate, benefit profoundly from standardization, bringing forth predictability and coherence, while automation paves the way for diminishing human-driven inconsistencies and bolstering deployment speeds. In containerized environments, these challenges are amplified due to the need to consistently deploy lightweight, encapsulated workloads across heterogeneous infrastructure with varying APIs and orchestration logic. Implications: For researchers, the quest for optimized frameworks championing standardization and automation within multi-cloud environment presents a promising research trajectory. For practitioners, adeptness with standardized methodologies and tools, combined with proficiency in automation implementation, becomes crucial in navigating the multi-cloud management maze efficiently
@@
Security and Compliance in multi-cloud environment: Security remains paramount in cloud deployments. Fernandes et al. underscore that the labyrinth of diverse platforms and differing standards in multi-cloud configurations amplifies security complexities [176]. Challenges articulated as Security Deployment in Multi-container Environment and Malicious Service Management Deployment in Table LABEL:tab:Deployment-solutions accentuate the primacy of security considerations. Propounded solutions like Cross-Container Isolation and Collusion-Resilient Trust Aggregation Technique spotlight the industry’s shift towards advanced and meticulous security protocols within multi-cloud paradigms. The ever-evolving landscape of multi-cloud deployments demands security solutions that are not only rigorous but also malleable and formidable. To attain comprehensive security across diverse platforms, strategies must encapsulate both depth of understanding and breadth of application. Containerization introduces new layers of isolation and deployment granularity, which—when combined with multi-cloud distribution—creates a unique surface of vulnerability that necessitates tailored controls at both the container and orchestration levels. Implications: For the academic community, pioneering avant-garde security solutions apt for multi-cloud environment offers a fertile domain of inquiry. Meanwhile, practitioners should enshrine principles of data integrity, trust assurance, and vulnerability mitigation at the heart of their multi-cloud deployment blackprints.
@@
Integration, Monitoring, and Infrastructure Management: As highlighted by Ferrer et al. [177] the ability to operate seamlessly across disparate cloud providers and maintain flexible infrastructures is paramount in harnessing the full potential of multi-cloud deployments. The presented challenges, including Network Latency Deployment and Vendor Lock-in Prevention Deployment, underscore the indispensability of agility and cross-platform compatibility. Proposed solutions such as VM Selection for Composite Applications and Interoperability Layer Above Cloud Infrastructure accentuate the importance of crafting adaptable infrastructure and formulating strategies to mitigate vendor entrenchment. Here, containerization plays a distinct role by enabling application components to be packaged once and executed across diverse cloud platforms, yet this same portability necessitates careful orchestration, consistent networking, and compatibility layers across varied cloud APIs and runtime environments. In the context of multi-cloud environment, flexibility and interoperability transcend mere advantageous attributes; they become pivotal determinants shaping adaptability, scalability, and the sustained success of deployment blackprints. Implications: For the academic community, delving into novel approaches that champion infrastructure pliability and seamless operation across diverse cloud platforms emerges as a prospective research avenue. For industry professionals, anchoring strategies around flexibility and interoperability not only optimizes current deployments but also fortifies them against future technological evolutions and shifting business dynamics.
4.7 Monitoring Challenges and Solutions (RQ4 - RQ5)
@@
Enhancing Performance Stability through Multi-Cloud Monitoring: In the context of multi-cloud environment, ensuring consistent and reliable performance monitoring for containerized applications can be challenging due to the dynamic nature of cloud resources. Existing studies (e.g., [168] [178]) highlighted the impact of varying resource availability on the performance of containerized applications across multiple clouds. They emphasized the need for performance benchmarking to track fluctuations and maintain a consistent user experience. The solutions presented in our framework, such as benchmarking performance variation and utilizing Kubernetes for monitoring, align with existing studies suggestion for benchmarking. Unlike traditional single-cloud deployments, containerized applications spanning multiple clouds must be monitored for platform-specific latency, orchestration timing mismatches, and fragmented telemetry data pipelines. Our proposed framework expands on this by incorporating Kubernetes for improved multi-cloud performance monitoring. Kubernetes, while originally designed for single-cluster orchestration, plays a central role in aggregating monitoring data across containerized services deployed in heterogeneous clouds, helping unify performance visibility. Our proposed framework expands on this by incorporating Kubernetes for improved multi-cloud performance monitoring. Our framework also suggest reliable performance monitoring requires a combination of standardized benchmarking techniques and cloud-specific tools like Kubernetes. This approach can effectively address performance variations and ensure a consistent user experience across diverse cloud environments. Implications: For practitioners, adopt benchmarking practices to track performance fluctuations, and use Kubernetes for unified multi-cloud performance monitoring. For researchers, explore further refinements in benchmarking methodologies and investigate the impact of dynamic multi-cloud environment on performance consistency.
@@
Streamlining Resource and Network Management across Multi-Clouds: Managing network communications, system images, and resource utilization in a multi-cloud setting poses challenges due to the diversity of cloud platforms. Sedghpour et al. discussed the complexities of managing network communications in multi-cloud environment. They emphasized the need for standardized approaches like service mesh to enhance network visibility and control [179]. Solutions in our proposed framework, such as using Docker Server for containerized services and Prometheus/Grafana for resource monitoring, align with Sedghpour et al.’s [179] suggestion for standardized approaches. However, container-based deployments compound the difficulty by adding layers of ephemeral, distributed services that must be monitored across volatile container lifecycles. This increases the burden of synchronization and observability. Additionally, our framework extends these solutions to address broader resource management challenges. Containers, especially when orchestrated over multiple cloud platforms, generate high-churn environments that require lightweight, highly adaptive monitoring strategies to ensure timely insights and cost control. Additionally, our framework extends these solutions to address broader resource management challenges. Employing standardized tools like Docker Server, along with platforms like Prometheus and Grafana, can help streamline resource management and network communication monitoring across diverse cloud platforms. Implications: For practitioners, implement Docker Server and utilize platforms like Prometheus and Grafana for efficient resource monitoring and management across multi-cloud environment. For researchers, investigate further integration possibilities for standardized resource management tools and explore the impact of such tools on overall system performance and reliability.
@@
Strengthening Security, Compliance, and Trust: Ensuring data privacy, continuous security control, and compliance in multi-cloud containerized applications is essential but challenging due to the distributed nature of cloud resources. Existing studies (e.g., [167] [6]) discussed the challenges of ensuring security and compliance across multi-cloud environment. They highlighted the importance of continuous monitoring and enforcement mechanisms to address security and compliance gaps. Solutions presented in our proposed framework, such as using machine learning for performance supervision and employing security assurance platforms, are aligned with existing studies that emphasis on continuous monitoring and enforcement for security and compliance. Containerized deployments introduce distinct security risks, e.g., image poisoning, privilege escalation between co-resident containers, and cross-platform trust boundary misconfigurations that demand container-aware monitoring solutions. Implementing advanced techniques like machine learning for monitoring and adopting security assurance platforms can significantly enhance security, compliance, and trust management in multi-cloud containerized applications. Implications: For practitioners, integrate machine learning-based monitoring and adopt security assurance platforms to ensure continuous security control and compliance across diverse cloud environments. For researchers, explore the scalability and effectiveness of machine learning-driven monitoring techniques and investigate novel methods for enforcing security and compliance in multi-cloud setups.
4.8 Tools and Frameworks (RQ6)
@@
Diverse Cloud Services by Leading Providers: Our study reveals a diverse array of cloud services, including IaaS, PaaS, and SaaS offerings from major providers like AWS, Azure, and Google Cloud, emphasizing the extensive options available to developers (see Table 15). This underscores the significant range of choices available to developers. Noteworthy cloud service providers highlighted in our study include Amazon EC2, Google Cloud App Engine, and Azure, further underscoring this finding. In multi-cloud environments, selecting services becomes more complex due to differing billing models, feature parity, and integration capabilities across providers—requiring tools that support cross-provider abstraction and orchestration. Implications: For practitioners, our outcomes underscore the critical necessity of meticulously evaluating and selecting the most fitting service model based on the unique demands of a given project. This strategic approach ensures the best alignment between technology and goals. Moreover, the implications for researchers are promising: an avenue emerges for in-depth exploration of trade-offs and the identification of optimal practices while navigating the intricate array of cloud service models. This exploration holds potential for enhancing both the efficiency and effectiveness of cloud solutions.
@@
Docker and Kubernetes: Pioneers of Cross-Cloud Containerization: Our investigation spotlights the profound significance of Docker and Kubernetes in streamlining the containerization and orchestration of applications across a spectrum of cloud platforms. Docker adeptly encapsulates applications and their dependencies within self-contained containers, while Kubernetes takes center stage as the orchestration juggernaut, automating deployment, scaling, and management of these containers. These technologies receive direct validation through table references, serving as prime exemplars. However, our analysis also reveals that deploying Kubernetes in a multi-cloud environment demands adaptations, such as federation control planes, multi-zone DNS, and workload placement strategies tailored to cross-cloud latencies and SLAs. Our research yields a compelling revelation: Docker and Kubernetes hold pivotal roles in ensuring consistent application deployment and seamless scalability, effectively simplifying the complexities of multi-cloud environment. Implications: Practitioners stand to gain significantly by adopting Docker and Kubernetes, as this potent combination bolsters efficiency and cultivates heightened portability. Researchers can explore emerging gaps in areas such as secure cross-cloud scheduling, fault tolerance across zones, and policy-compliant orchestration.
@@
The Rise of DevOps: Embracing Automation Tools: Our exploration spotlights the important role of DevOps practices and tools such as Puppet, Chef, and Terraform within contemporary software development landscapes. This emphasis revolves around the power of automation in facilitating streamlined deployment and management processes. These tools, by enabling consistent provisioning and deployment of infrastructure, effectively curtail errors and ensure the replicability of processes. The table substantiates this notion through references to Puppet, Chef, and Terraform. Within a multi-cloud context, these tools must be adapted to handle provider-specific APIs, authentication schemes, and compliance constraints, which introduces configuration drift and policy management complexity.. Implications: Practitioners are advised to seamlessly integrate DevOps principles into their workflows by harnessing the prowess of these tools. Researchers may investigate how DevOps pipelines can be extended to support containerized workloads that span multiple clouds, ensuring traceability, consistency, and auditability across heterogeneous execution environments.
5 Threats to Validity
@@
This systematic mapping study is susceptible to various threats that could influence its outcomes. To address these potential threats, we adhered to the established guidelines for conducting SMSs and SLRs as outlined in [17] and [180]. We further categorized and analyzed the validity threats to our study based on the four distinct types of validity threats mentioned in [181] and [182]. In the subsequent sections, we delve into the specific validity threats that pertain to the different phases of this SMS.
5.1 Internal Validity
@@
  • Study Selection: We have outlined the study screening and selection process in Table 3. To ensure objectivity and eliminate personal bias in study selection, we implemented a two-phase approach: (i) initial study screening, and (ii) qualitative evaluation of the shortlisted studies. During this procedure, the lead two authors conducted the study screening based on the criteria explicitly detailed in Section 3. To assess the objectivity of the screening process and the level of agreement between the two authors, we applied Cohen’s Kappa [22]. In cases where the lead authors did not agree, the second and third authors independently reviewed the disputed studies to reach a consensus. All the researchers involved in this study have extensive knowledge and research experience in containerization applications within a multi-cloud environment.
  • Data Extraction: Bias from researchers during data extraction can pose a significant challenge in both SMSs and SLRs. To address this potential issue, we developed a standardized data extraction form (see Table 5) to ensure consistent data retrieval. Initially, the first and fourth authors extracted the data. If any uncertainties arose regarding the extracted data, comprehensive discussions were convened among all authors. Following the recommendations in [182], a subset of the extracted data was cross-verified by the second and third authors.
  • Bias on Themes Classification: Misclassification of data and primary studies may result in biases stemming from subjective interpretation. To minimize this risk, we adhered to the thematic analysis guidelines established by Braun et al. [26], and implemented a six-step process for thematic classification, as detailed in Section 2.4.
  • Data Synthesis: We employed both qualitative and quantitative approaches to assess the gathered data. Potential biases in data synthesis could influence the interpretation of our findings. To address this concern, we used open coding and constant comparison techniques from Grounded Theory [27] to analyze the qualitative data extracted from the selected studies.
  • Transparency of Quality Assessment Process: While our replication package [11] provides a detailed breakdown of quality assessment scores for each study and each criterion, it does not include per-study justifications or annotations explaining how individual scores were assigned. This may present a limitation in terms of external reproducibility, as other researchers may not fully reconstruct the decision-making rationale behind each score. Although we followed a consistent and internally validated scoring logic (e.g., peer-reviewed venues scored higher, empirical data presence was required for full marks), we recognize that the absence of a fully documented scoring rubric or narrative rationale could reduce transparency. However, to partially mitigate this, we have explicitly described our scoring criteria and interpretation rules in Section 2.3.3. This clarification improves traceability by explaining the rationale used across all assessed studies. The limitation remains inherent to balancing the breadth and depth of quality assessment in a large-scale review and is acknowledged as a tradeoff in terms of replication of this study.
5.2 External Validity